A country has no travelers’ checks and it has the data given below. What is the value for

checking account balances? __________

Currency issued = 1400 Vault cash= 400 rrr = 10%

err = 5 % bank reserves with the Fed = 1600cr = ?

Total deposits in banks = 10,000 govt budget deficit = 300

Note: M1 ={ (1 + cr)/(rrr + err + cr)} x MB
Answer #1

Calculate the monetary base (MB) -

Monetary base = Currency issued + Bank reserves with the Fed + Valut cash

Monetary base = 1400 + 1600 + 400

Monetary base = 3400

The monetary base is 3400.

Calculate the currency deposit ratio =

cr = curreny/deposit = 1400/10000 = 0.14

The currency-deposit ratio is 0.14

Calculate the money multiplier -

Money multiplier = (1 + cr)/(rrr + err + cr)

Money multiplier = (1 + 0.14)/(0.10 + 0.05 + 0.14)

Money multiplier = 1.14/0.29 = 3.93

The money multiplier is 3.93

Calculate the M1 money supply -

M1 = Money multiplier * MB

M1 = 3.93 * 3400 = 13,362

The M1 money supply is 13362.

Calculate the checking account balances -

Checking account balances = M1 money supply - Currency issued

Checking account balances = 13362 - 1400 = 11962

The checking account balances is 11,962